Two numbers differ by 57. Their sum is 185. Find the numbers.
Answer by benazir.sj@gmail.com(109)

let the numbers be x and y
sum of numbers is 185
x+y =185-----------(1)
differnce is 57
x-y =57------------- (2)
(1) +(2) =
x+y=185
x-y=57
=> 2x= 242
x=242/2
x=121
x-y=57
121 -y=57
-y=57-121
-y= -64
y=64
therfore the two numbers are
121 and 64